INVESTIGATION OF THE COLLISION-INDUCED-ABSORPTION CONTRIBUTION TO THE 1.27μm a^{1}Δ_{g} - X^{3}Σ_{g}^{−} BAND OF O_{2}

C.L. LUGEZ, W.J. LAFFERTY, AND \underline{G.T. FRASER}, OPTICAL TECHNOLOGY DIVISION, NA-TIONAL INSTITUTE OF STANDARDS AND TECHNOLOGY, GAITHERSBURG, MD 20899.

The collision-induced absorption structure accompanying the magnetic-dipole-allowed 1.27 μm a^{1}Δ_{g} - X^{3}Σ_{g}^{−} band of O_{2}, originally observed in the laboratory by Badger, Wright, and Whitlock, has been measured at 0.03 cm^{−1} resolution using a Fourier-transform infrared spectrometer. Spectra were recorded at an optical pathlength of 84 m for pure O_{2}, at temperatures between 253 K to 293 K and pressures to 9 atm. Several spectra were also measured for room-temperature O_{2}/N_{2} mixtures. These results presently being analyzed will allow comparison with recent measurements of these bands in the atmosphere by Mlawer, Clough, Brown, Stephen, Landry, Goldman, and Murcray.